40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 236 of 341  Not logged in ELOG logo
ID Date Author Type Category Subject
  5305   Thu Aug 25 17:57:35 2011 SureshUpdateSUSBroken UL magnet on ITMX

Quote:

Dmass just reminded me that the usual procedure is to bake the optics after the last gluing, before putting them into the chambers.  Does anyone have opinions on this? 

On the one hand, it's probably safer to do a vacuum bake, just to be sure.  On the other hand, even if we could use one of the ovens immediately, it's a 48 hour bake, plus cool down time.  But they're working on aLIGO cables, and might not have an oven for us for a while.  Thoughts?

I think we should follow the established procedure in full, even though it will cost us a few more days.  I dont think we should consider the vacuum bake as something "optional".  If the glue has any volatile components they could be deposited on the optic resulting in a change in the coating and consequently optical loss in the arm cavity.

 

 

  5304   Thu Aug 25 17:40:07 2011 DmassUpdateComputer Scripts / Programselog broke, fixed

elog died b/c someone somewhere did something which may or may not have been innocuous. I ran the script in /cvs/cds/caltech/elog to restart the elog (thrice).

 

I have now banned Warren from clicking on the elog from home

  5303   Thu Aug 25 17:14:49 2011 steveUpdateVACclean room fashion changes

Jamie is modeling our next generation  in-vac & clean room bonny suit that Jenne and myself already tested.

It is quite bearable with our traditional cleanroom beret-bouffant cap. Please use these in the future.

This will help to avoid the farther degradation of somewhat dusty 40m vac envelope.

It is the required dress code to enter the clean assembly room in the 40m.

 We have small, med, large and x-large in stock. I'm getting larger sizes.

It will not allow certain people to climb inside the vacuum chamber in dirty pants.

Attachment 1: P1080185.JPG
P1080185.JPG
Attachment 2: P1080183.JPG
P1080183.JPG
  5302   Thu Aug 25 15:20:03 2011 JenneUpdateSUSBroken UL magnet on ITMX

Dmass just reminded me that the usual procedure is to bake the optics after the last gluing, before putting them into the chambers.  Does anyone have opinions on this? 

On the one hand, it's probably safer to do a vacuum bake, just to be sure.  On the other hand, even if we could use one of the ovens immediately, it's a 48 hour bake, plus cool down time.  But they're working on aLIGO cables, and might not have an oven for us for a while.  Thoughts?

  5301   Thu Aug 25 13:10:42 2011 JenneUpdateSUSDrag wiping

As we have seen in the past, both of the ITMs were more dusty than the ETMs, presumably because we have the vertex open much more often than the ends.  Kiwamu and I wiped all of the optics until we could no longer see any dust particles within a ~1.5 inch diameter area around the center. 

Since we have ITMX out for magnet gluing, I'll probably drag wipe both front and back surfaces before putting it back in the suspension cage.  All of the optics have clear dust on the AR surfaces, but we can't get to that surface while the optics are suspended.  For the ETMs this isn't too big of a deal, but it does concern me a bit for the ITMs and other transmissive optics we have.  I don't think it's bad enough yet though to warrant removing optics from suspensions just to wipe them.

  5300   Thu Aug 25 08:12:09 2011 steveUpdateSUSETMY & ITMY wiped, replaced, vac door on

Quote:

We've closed up ETMX:

  • the optic was drag wiped
  • the suspension tower was put back in place
  • earthquake stops were backed off the appropriate number of turns, and de-ionized
  • chamber door was put on

 jamie, jenne, kiwamu, suresh, steve

 

ETMY and ITMY were treated the same way as ETMX. The BS chamber was closed with heavy vac door yesterday also. The IOO access connector's inner jamnuts are torqued to 45 ft/lbs as all vac door bolts.

The vac envelope is ready for pumpdown condition, except ITMX chamber with light atm door cover.

Jenne will summeries the condition of dust on the  TMs before and after the drag wipes.

  5299   Wed Aug 24 17:05:11 2011 JenneUpdateSUSBroken UL magnet on ITMX

Quote:

The ITMX tower was shipped into the Bob's clean room to put the magnet back on. 

 Repair work is delayed.  I need the "pickle pickers" that hold the magnet+dumbbell in the gluing fixture, for gluing them to the optic.  Here at the 40m we have a full set of SOS gluing supplies, except for pickle pickers.  We had borrowed Betsy's from Hanford for about a year, but a few months ago I returned all of the supplies we had borrowed.  Betsy said she would find them in her lab, and overnight them to us.  Since the problem occurred so late in the day, they won't get shipped until tomorrow (Thursday), and won't arrive until Friday.

I also can't find our magnet-to-dumbbell gluing fixture, so I asked her to send us her one of those, as well. 

I have 2 options for fixing ITMX.  I'll write down the pros and cons for each, and we can make a decision over the next ~36 hours.

OPTIONS:

(#1) Remove dumbbell from optic.  Reglue magnet to dumbbell. Reglue magnet+dumbbell to optic.

(#2) Carefully clean dumbbell and magnet, without breaking dumbbell off of optic.  Glue magnet to dumbbell.

PROS:

(#1) Guarantee that magnet and dumbbell are axially aligned.

(#2) Takes only 1 day of glue curing time.

CONS:

(#1) Takes 2 days of glue curing time. (one for magnet to dumbbell, one for set to optic.)

(#2) Could have slight mismatch in axis of dumbbell and magnet.  Could accidentally drop a bit of acetone onto dumbbell-to-optic glue, which forces us into option 1, since this might destroy the integrity of the glue joint (this would take only the 2 days already required for option 1, it wouldn't force us to take 2+1=3 days).

  5298   Wed Aug 24 16:13:36 2011 kiwamuUpdateSUSbroke UL magnet on ITMX

I broke the UL magnet on ITMX

The ITMX tower was shipped into the Bob's clean room to put the magnet back on.

 

 Since we found that all the magnets were relatively high (#5296) in the shadow sensors, we decided to slide the OSEM holder bar upward.

During the work, I haven't made the OSEMs far enough from the magnets.

So the magnets and OSEMs touched as I moved the holder.

Then the UL magnets were broken off and fell into the UL coil.

 

  5297   Wed Aug 24 12:08:56 2011 jamieUpdateSUSITMX, ETMX, ETMY free swinging

ITMX: 998245556

ETMX, ETMY: 998248032

  5296   Wed Aug 24 11:40:21 2011 jamie, jenne, kiwamu, suresh, steveUpdateSUSproblem with ITMX

ITMX was drag wiped, and the suspension was put back into place.  However, after removing all of the earthquake stops we found that the suspension was hanging in a very strange way.

The optic appears to heavily pitched forward in the suspension.  All of the rear face magnets are high in their OSEMs, while the SIDE OSEM appears fine.  When first inspected, some of the magnets appeared to be stuck to their top OSEM plates, which was definitely causing it to pitch forward severely.  After gently touching the top of the optic I could get the magnets to sit in a more reasonable position in the OSEMs.  However, they still seem to be sitting a little high.  All of the PDMon values are also too low:

  nominal now
UL 1.045 0.767
UR 0.855 0.718
LR

0.745

0.420

LL

0.780

0.415
SD

0.840

0.752

Taking a free swing measurement now.

  5295   Wed Aug 24 11:30:27 2011 jamie, jenne, kiwamu, suresh, steveUpdateSUSETMX wiped, replaced, door on

We've closed up ETMX:

  • the optic was drag wiped
  • the suspension tower was put back in place
  • earthquake stops were backed off the appropriate number of turns, and de-ionized
  • chamber door was put on
  5294   Wed Aug 24 09:11:19 2011 jamieUpdateSUSETMY SUS update: looks good. WE'RE READY TO CLOSE

We ran one more free swing test on ETMY last night, after the last bit of tweaking on the SIDE OSEM.  It now looks pretty good:

ETMY ETMY.png       pit     yaw     pos     side    butt
UL   -0.323   1.274   1.459  -0.019   0.932 
UR    1.013  -0.726   1.410  -0.050  -1.099 
LR   -0.664  -1.353   0.541  -0.036   0.750 
LL   -2.000   0.647   0.590  -0.004  -1.219 
SD    0.021  -0.035   1.174   1.000   0.137 
4.23371

 So I declare: WE'RE NOW READY TO CLOSE UP.

  5293   Tue Aug 23 18:25:56 2011 jamieUpdateSUSSRM diagnalization OK

By looking at a longer data stretch for the SRM (6 hours instead of just one), we were able to get enough extra resolution to make fits to the very close POS and SIDE peaks.  This allowed us to do the matrix inversion.  The result is that SRM looks pretty good, and agrees with what was measured previously:

SRM SRM.png        pit     yaw     pos     side    butt
UL    0.869   0.975   1.140  -0.253   1.085  
UR    1.028  -1.025   1.083  -0.128  -1.063  
LR   -0.972  -0.993   0.860  -0.080   0.834  
LL   -1.131   1.007   0.917  -0.205  -1.018  
SD    0.106   0.064   3.188   1.000  -0.011 
4.24889

 

  5292   Tue Aug 23 17:51:37 2011 KeikoUpdateLSCTolerance of PRC, SRC, MICH length = 2 mm ?

Keiko, Kiwamu

We noticed that we have used wrong code for MICH degree of freedom for both of the ELOG entries on this topic (cavity lengths tolerance search). It will be modified and posted soon.

  5291   Tue Aug 23 17:45:22 2011 jamieUpdateSUSITMX, ITMY, ETMX clamped and moved to edge of tables

In preparation for tomorrow's drag wiping and door closing, I have clamped ITMX, ITMY, and ETMX with their earthquake stops and moved the suspension cages to the door-edge of their respective tables.  They will remain clamped through drag wiping.

ETMY was left free-swinging, so we will clamp and move it directly prior to drag wiping tomorrow morning.

  5290   Tue Aug 23 17:21:45 2011 kiwamuUpdateSUSfree swinging test for ETMY

 

Excited ETMY

Tue Aug 23 17:20:45 PDT 2011
998180460

 

  5289   Tue Aug 23 16:23:33 2011 JenneUpdateVACAccess connector in place

[Steve, Bob, Jamie, Kiwamu, Valera, Jenne]

The access connector is now in place, in preparation for pump-down.  Tomorrow (hopefully) we will do all the other doors.

 

  5288   Tue Aug 23 14:49:14 2011 jamie, jenne, kiwamu, suresh, keikoUpdateSUSAdjustment of ETMY, issue with ITMY whitening

Before lunch we took a closer look at two of the suspensions that were most problematic: ITMY and ETMY.  Over lunch we took new free swinging data.  Results below:

  • For ITMY we discovered that the whitening on the UL sensor was not switching.  This was causing the UL sensor to have a different response, with a steeper roll of, which was causing all of the transfer function estimates to the other sensors to have large imaginary components.   We took new free swing data with all of the whitening turned OFF.  The result is a much improved matrix and diagnalization.  The input matrix elements are mostly the same, but the coupling is basically gone.  We'll fix the whitening after the pump down.
ITMY ITMY.png       pit     yaw     pos     side    butt
UL    0.157   1.311   1.213  -0.090   0.956 
UR    1.749  -0.490   0.886  -0.038  -1.042 
LR   -0.251  -2.000   0.787  -0.007   1.066 
LL   -1.843  -0.199   1.114  -0.059  -0.936 
SD   -0.973  -0.205   1.428   1.000   0.239 
4.34779
  • ETMY has a very problematic SIDE OSEM.  The magnet does not line up with the OSEM axis, and since there is no lateral adjustment in the side OSEMs, there's not much we can do about this.  We're using aluminum foil to wedge the OSEM over as far as possible, but it's not quite enough.  With the OSEM plates horizontal there is a lot of POS->SIDE coupling.  With the OSEM plates vertical, the magnetic sits a little too close to the rear face, which can cause the magnet to get stuck to the LED plate.  We're trying to decide where to leave it now, but the new diagnalization with the OSEM plates vertical is definitely better: 
ETMX ETMY.png        pit     yaw     pos     side    butt
UL   -0.138   1.224   1.463  -0.086   0.944  
UR    0.867  -0.776   1.501  -0.072  -1.051  
LR   -0.995  -0.896   0.537  -0.045   0.754  
LL   -2.000   1.104   0.499  -0.059  -1.251  
SD    0.011   0.220   1.917   1.000   0.224 
 4.42482
  5287   Tue Aug 23 11:57:22 2011 kiwamuUpdateSUSfree swinging test during lunch time

excited all the optics. (with ITMY WTF OFF)

Tue Aug 23 11:52:52 PDT 2011
998160788

  5286   Tue Aug 23 10:38:27 2011 jamieUpdateSUSSUS update

SUS update before closing up:

  • MC1, MC2, ITMX look good
  • MC3, PRM look ok
  • SRM pos and side peaks are too close together to distinguish, so the matrix is not diagnalizable.  I think with more data it should be ok, though.
  • all ITMY elements have imaginary components
  • ITMY, ETMX, ETMY appear to have modest that swapped position:
    • ITMY: pit/yaw
    • ETMX: yaw/side
    • ETMY: pos/side
  • MC3, ETMX, ETMY have some very large/small elements

Not particularly good.  We're going to work on ETMY at least, since that one is clearly bad.

OPTIC   M cond(B)
MC1 MC1.png       pit     yaw     pos     side    butt
UL    0.733   1.198   1.168   0.050   1.057 
UR    1.165  -0.802   0.896   0.015  -0.925 
LR   -0.835  -1.278   0.832  -0.002   0.954 
LL   -1.267   0.722   1.104   0.032  -1.064 
SD    0.115   0.153  -0.436   1.000  -0.044
4.02107
MC2 MC2.png        pit     yaw     pos     side    butt
UL    1.051   0.765   1.027   0.128   0.952  
UR    0.641  -1.235   1.089  -0.089  -0.942  
LR   -1.359  -0.677   0.973  -0.097   1.011  
LL   -0.949   1.323   0.911   0.121  -1.096  
SD   -0.091  -0.147  -0.792   1.000  -0.066 
4.02254
MC3  MC3.png        pit     yaw     pos     side    butt
UL    1.589   0.353   1.148   0.170   1.099  
UR    0.039  -1.647   1.145   0.207  -1.010  
LR   -1.961  -0.000   0.852   0.113   0.896  
LL   -0.411   2.000   0.855   0.076  -0.994  
SD   -0.418   0.396  -1.624   1.000   0.019
3.60876
PRM  PRM.png        pit     yaw     pos     side    butt
UL    0.532   1.424   1.808  -0.334   0.839  
UR    1.355  -0.576   0.546  -0.052  -0.890  
LR   -0.645  -0.979   0.192   0.015   0.881  
LL   -1.468   1.021   1.454  -0.267  -1.391  
SD    0.679  -0.546  -0.674   1.000   0.590 
 5.54281
BS  BS.png        pit     yaw     pos     side    butt
UL    1.596   0.666   0.416   0.277   1.037  
UR    0.201  -1.334   1.679  -0.047  -0.934  
LR   -1.799  -0.203   1.584  -0.077   0.952  
LL   -0.404   1.797   0.321   0.247  -1.077  
SD    0.711   0.301  -3.397   1.000   0.034 
 5.46234
SRM  NA  NA  NA
ITMX  ITMX.png        pit     yaw     pos     side    butt
UL    0.458   1.025   1.060  -0.065   0.753  
UR    0.849  -0.975   1.152  -0.199  -0.978  
LR   -1.151  -1.245   0.940  -0.243   1.217  
LL   -1.542   0.755   0.848  -0.109  -1.052  
SD   -0.501  -0.719   2.278   1.000  -0.153
 4.4212
ITMY  ITMY.png        pit     yaw     pos     side    butt
UL    0.164   1.320   1.218  -0.086   0.963  
UR    1.748  -0.497   0.889  -0.034  -1.043  
LR   -0.252  -2.000   0.782  -0.005   1.066  
LL   -1.836  -0.183   1.111  -0.058  -0.929  
SD   -0.961  -0.194   1.385   1.000   0.239 
 4.33051
ETMX ETMX.png        pit     yaw     pos     side    butt
UL    0.623   1.552   1.596  -0.033   1.027  
UR    0.194  -0.448   1.841   0.491  -1.170  
LR   -1.806  -0.478   0.404   0.520   0.943  
LL   -1.377   1.522   0.159  -0.005  -0.860  
SD    1.425   3.638  -0.762   1.000  -0.132 
 4.89418
ETMY ETMY.png        pit     yaw     pos     side    butt
UL    0.856   0.007   1.799   0.241   1.005  
UR   -0.082  -1.914  -0.201  -0.352  -1.128  
LR   -2.000   0.079  -0.104  -0.162   0.748  
LL   -1.063   2.000   1.896   0.432  -1.119  
SD   -0.491  -1.546   2.926   1.000   0.169 
 9.11516

 

  5285   Tue Aug 23 09:40:37 2011 kiwamuUpdateGeneralRe: AS clipping fixed

Indeed it was suspenseful.

We tried finding where the clipping happened, but we couldn't find any obvious clippings.

So we checked centering of the beams on all the optics associated with the AS path, starting from BS, SR3,... to the AS optical bench.

And during the work some of them were recentered.

At the end we found no clipping. To make sure we tested the available range (no clipping range) by exciting the angular motion of BS with AWG (f ~ 1Hz, a ~ 1000).

The beam looked successfully coming out at the most of the angular oscillation point.

Quote from #5284

Where was the AS clipping?! Ah, the suspense...

 

  5284   Tue Aug 23 06:49:24 2011 AnamariaUpdateGeneralmore in-vac work : AS clipping fixed and OSEM/oplev adjustment

Where was the AS clipping?! Ah, the suspense...

Quote:

  + fixed the AS clipping issue

 

Quote from #5275

We need to check/fix the AS beam clipping and once it's done we will readjust the OSEM mid range and the oplevs.

 

 

 

  5283   Tue Aug 23 02:03:04 2011 SureshUpdateIOOMC realigned and spot positions recentered

After the MC1 and MC3 OSEMs were  repositioned  MC had to be realigned and the beam spots had to be recentered on the actuation nodes.  

To do that I had to change the input beam direction into the MC  and the coil offsets.   

I also measured the resultant spot positions

spot positions in mm (MC1,2,3 pit MC1,2,3 yaw):
    0.1354   -0.2522   -0.1383   -1.0893    0.7122   -1.5587

mcdecenter20110822.png

 

The MC1 and MC3 yaw can be improved further after the chambers are closed and evacuated.  The PZT adjustments needed to realign the input beam pointing are quite small and should not pose a problem.

 

 

  5282   Tue Aug 23 01:09:44 2011 kiwamuUpdateSUSfree swinging test

excited all the optics ---

Tue Aug 23 01:08:00 PDT 2011
998122096

  5281   Tue Aug 23 01:05:40 2011 JenneUpdateTreasureAll Hands on Deck, 9am!

We will begin drag wiping and putting on doors at 9am tomorrow (Tuesday). 

We need to get started on time so that we can finish at least the 4 test masses before lunch (if possible). 

We will have a ~2 hour break for LIGOX + Valera's talk.

 

I propose the following teams:

(Team 1: 2 people, one clean, one dirty) Open light doors, clamp EQ stops, move optic close to door.  ETMX, ITMX, ITMY, ETMY

(Team 2: K&J) Drag wipe optic, and put back against rails. Follow Team 1 around.

(Team 3 = Team 1, redux: 2 people, one clean, one dirty) Put earthquake stops at correct 2mm distance. Follow Team 2 around.

(Team 4: 3 people, Steve + 2) Close doors.  Follow Team 3 around.

Later, we'll do BS door and Access Connector.  BS, SRM, PRM already have the EQ stops at proper distances.

 

  5280   Tue Aug 23 00:55:13 2011 JenneUpdateGeneralIFO ready for doors

[Kiwamu, Jenne]

After the IFO was aligned in air one final time, we tapped on a few OSEMs until we were happy with all of the centering of all of the optics' OSEMs.  All are within 0.05 of their halfway values, with the exception of one each on MC1 and MC3, one of which is within 0.06, and the other 0.08.  Because of the realignment pain of dealing with MC OSEMs, we elected to leave these alone.  Also, since we obviously didn't open the MC2 tank, we don't know how they are, although the numbers look reasonable. 

Also, we took photos (to be posted on Picasa in a day or two) of all the main IFO magnet-in-OSEM centering, as best we could.  SRM, BS, PRM all caused trouble, due to their tight optical layouts.  We got what we could.  Various people have been looking at these for the past 2 weeks, and I think they're all fine, even if we didn't get stellar photos.

We are now prepared for pumping.  For real this time.

 

  5279   Mon Aug 22 21:32:10 2011 kiwamuUpdateGeneralmore in-vac work : AS clipping fixed and OSEM/oplev adjustment

[Keiko / Jenne / Jamie / Kiwamu]

 We did the following things today :

  + fixed the AS clipping issue

  + realigned all the oplevs

  + checked and adjusted the all OSEM DC values, including PRM, SRM, BS, ITMs, ETMs, MC1 and MC3

 

Since we touched the OSEMs the alignment has changed somewhat.

Right now Jenne, Suresh and I are working on the "confirmation alignment".

Once we find the alignment is still good (steerable by the PZTs and the DC coil bias), tomorrow we will do the drag&wipe and door closing.

Quote from #5275

We need to check/fix the AS beam clipping and once it's done we will readjust the OSEM mid range and the oplevs.

 

  5278   Mon Aug 22 20:37:43 2011 AnamariaConfigurationRF SystemPlan for install of 3f PDs

I made a quick sketch of how to include two more RF PDs on the REFL beam, given the space we have on the table. We want to install REFL33 and REFL165, 3f signals for the the two modulation frequencies we are using. The point is to make the distance from first beam splitter the same to all PDs so that we can use only one lens before this BS to make the beam the right size. Currently there are 2 PDs on the refl beam, REFL11 and REFL55, predictably. So the drawing shows 4 PDs. Drawing is to scale but is a bit coarse. Hopefully we'll take pictures once we're done.

Reference from current BS splitting beam to the existing PDs.

Attachment 1: 40m3fReflPdLayout.pdf
40m3fReflPdLayout.pdf
  5277   Mon Aug 22 16:28:44 2011 SureshUpdateGeneralSRM tower shifted on Friday

[Kiwamu, Suresh]   This is a belated elog entry from last Friday night + Saturday morning! 

    We shifted the SRM tower across the beam and away from the door by about 5mm. 

                  After the input beam from MC was aligned to the Y-arm,  Kiwamu noticed that the AS beam was being clipped and that the correction had to start from SRM onwards as the beam had become offcentered on the SRM.  So we shifted the SRM tower by about 5mm away from the door and transverse to the beam and rotated it by a few degrees to center the OSEM offsets.  After this we aligned all optics along the AS beam path to extract the AS beam from the chamber.    We then aligned each optic in the vertex so that their beams overlap at the AS port with the reflection from ITMY.  Then we aligned BS to center the beam on ETMX and then looked for flashes from the Y arm.

      At this point Kiwamu checked and found that the input beam from the MC had shifted.  It was landing on the ITMY about 5mm below the center.  And (inexplicably) it was still centered on ETMY!  The Y- green which traced the cavity axis (since this was still flashing) was not coincident with the IR beam.   So all the work we did in aligning the AS beam and the vertex optics work was lost..... and had to be done again.

 

 

  5276   Mon Aug 22 11:40:08 2011 steveUpdateVACcranes checked

Cranes are  checked and they are ready for lifting. At the east end will use the manual Genei-lift to put door on.

  5275   Mon Aug 22 02:43:28 2011 kiwamuUpdateGeneralin-vac work

[Kieko / Kiwamu]

 The AS beam shows a little bit of clipping and right now this is the only concern for the alignment of the interferometer.

Other than that everything is okay including :

  + Beam centering on all the suspended optics

  + Arm cavity alignments. The Fabry-Perot fringes from both arm cavity were found on the AS camera.

  + POX/POY/POP, they are still successfully coming out from the chambers

  + IPPOS/ANG

  + Alignment of the green beams and the associated optics. Both green are reaching to the PSL table

 

We need to check/fix the AS beam clipping and once it's done we will readjust the OSEM mid range and the oplevs.

Then it will be ready for the drag/wipe and door closing.

  5274   Sat Aug 20 23:01:39 2011 JennyUpdatePSLTaking temperature step-response data: successes and tribulations

After finishing my last elog entry, I monitored the digital loop's error signal (the control signal for the fast loop) and the output to the laser heater remotely, (from West Bridge), using dataviewer. The ref cav surrounding can temperature was set to 36 degrees C.

With the loop closed and a gain of 0.008, after seeing the output voltage to the laser heater (TMP_OUTPUT) remain fairly constant and the error signal (TMP_INMON) stay close to zero for ~3 hours, I tried to step the temperature. (This was at 2am last night). I was working remotely from West Bridge. To step the temperature I used the following command:

ezcawrite C1:PSL-FSS_RCPID_SETPOINT 35.5

 

Rather than change the can temperature to 35.5 C, it outputted:

C1:PSL-FSS_RCPID_SETPOINT=0.

 

It had set the setpoint to 0 degrees C, which was essentially turning the heater off. I tried resetting it back to 36 and had no luck. I tried changing the syntax slightly.: ezcawrite C1:PSL-FSS_RCPID_SETPOINT=36 and ezcawrite C1:PSL-FSS_RCPID_SETPOINT (36). No success.

I ran over to the 40m and changed the temperature back to 36 manually. The in-loop temp sensor had decreased to 31.5 degrees C before I was able to step the setpoint back up. The system seems to have recovered from this large impulse though, and the laser has remained locked.

5hourwbigimpulse.jpg

 5hourwbigimpulse2.jpg

(5 hours of minute-trend data)

From left to right: 

Top: Out-of-loop can temp sensor; Voltage sent to heat can

Middle: signal sent to heat the laser (TMP_OUTPUT); room temp

Bottom: Error signal for slow loop (sampled control signal from fast loop); In-loop can temp sensor

 

At 9:30 this morning (7 and a half hours after accidentally setting the setpoint to zero), I came in to the 40m. TMP_OUTPUT was still decreasing but was slowing somewhat, so I decided to step the can temperature up half a decree to 36.5 C.

TMP_OUTPUT responded to the step, but it is also oscillating slowly with room-temperature changes, and these oscillations are on the same order as the step response. The oscillations look like the room-temp oscilations, but inverted. (TMP_OUTPUT reaches maxima when RMTEMP reaches minima). Oddly, there does not appear to be much of a time delay between the room temperature and TMP_OUTPUT signals. I would expect a time delay since there's a time constant for a room-temperature change to propagate into the cavity. Perhaps the laser itself is susceptible to room-temperature changes and those propagate into the laser cavity on a much faster time scale. I don't know the thermal coupling of ambient temperature changes into the laser.

23hoursbefore920pm.jpg

23hoursbefore920pm2.jpg

(24-hours of second-trend data)

 

Options are:

--If the system can handle it, do a larger temperature step (3 degrees, say), so that I can more clearly distinguish the oscillations with room temp from the step response.

--Insulate the cavity with foam (will in principle make the temperature over the can surrounding the ref cav more uniform and less affected by room temperature changes).

--Insulate the laser? Is this possible?

--Leave the system as is and, as a first approximation, fit the room-temp data to a sine wave and subtract it off somehow from my data to just see the step response.

--Don't bother with steps and just try to get the transfer function from out-of-loop temperature (RCTEMP, which is affected by temperature noise from the room) to TMP_OUTPUT via taking the Fourier transforms of both signals.

 

I'm flying out tomorrow morning, so I'll either need to figure out how to step the temperature set point of the can remotely, successfully, or I'll need someone to manually enter in the temperature steps for me in the control room.

  5273   Sat Aug 20 00:42:22 2011 KeikoUpdateLSCTolerance of PRC, SRC, MICH length = 2 mm ?

 Keiko, Kiwamu

 I have run Kiwamu's length tolerance code (in CVS iscmodeling, ArmTolerance.m & analyseArmTorelance.m ) for the vertex ifo.

In his previous post, he monte-carlo-ed the arm lengths and saw the histogram of the sensing matrix and the demodulation phase between POP55 MICH and POP55 SRCL. From these plots, he roughly estimated that the tolerance is about 1 cm (sigma of the rondom gaussian) and in that case POP55 MICH and SRCL is separated by the demodulation phase 60-150 degrees.

This time I put the length displacements of random gaussian on PRC, SRC, MICH lengths at the same time (Fig.1).

 

fig3.png

Fig. 1. History of random walk in PRC, SRC, MICH lengths parameter space. Same as Kiwamu's previous post, The position of the three degrees are randomly chosen with a Gaussian distribution function in every simulaton. This example was generated when \sigma = 1 cm for all the three lengths, where \sigma is the standard deviation of  the Gaussian function. The number of simulation is 1000 times.

When the sigma is 1 cm, we found that the sensing matrix is quite bad if you look at Fig. 2. In Fig.2 row POP55, although the desired degrees of freedoms are MICH and SRCL, they have quite a bit of variety. Their separation in the demodulation phase is plotted in Fig.3. The separation in the demodulation phase varies from 40 degrees to 140 degrees, and around 270 degrees. It is not good as ideally we want it to be 90.

drawing.pdf

Fig. 2 Histgram of the sensing signal power in the matrix when 1 cm sigma rondom gaussian is applied on PRC, SRC, MICH lengths. x axis it the signal power in log10.

 

 

fig4.png

  Fig.3 POP55 MICH and POP55 SRCL separation with the displacement sigma 1 cm. 

  

 Kiwamu suspected that PRC length as more strict tolerance than other two (SRC, MICH) for POP55, as 55MHz is fast and can be sensitive to the arm length change. So I ran the same monte-carlo with SRC, MICH displacements but no PRC displacements when sigma is the same, 1cm. The results were almost same as above, nothing obvious difference.

 

With 2mm sigma, the signal power matrix and the POP55 MICH and POP55 SRCL separation in the demodulation phase look good (Fig. 4 and Fig. 5). 

fig1.pdf 

 Fig.4 Signal power matrix when PRC, SRC, MICH lengths fractuate with random gaussian distribution with 2mm sigma. The signal powers are shown in log10 in x axis, and they do not vary very much in this case.

fig4.png 

 Fig.5 POP55 MICH and POP55 SRCL separation with the displacement sigma 2 mm. The separation of the two signal is 60-90 degrees, much better than when sigma is 1 cm. We may need to check 60 degree separation is really ok or not.

 

PRC SRC MICH lengths tolerances of 2 mm in the real world will be very difficult ! 

Next I will check what happens on 3f signals.

 

 

Quote:

 Required arm length = 37.7974 +/- 0.02 [m]

This is a preliminary result of the estimation of the Arm length tolerance.

This number was obtained from a simulation based on Optickle.
Note that the simulation was done by considering misplacements in only the arm lengths while keeping PRCL, SRCL and MICH at the ideal lengths.
Therefore the tolerance will be somewhat tighter if misplacements in the central part are taken into account.

Next : check 3f signals, and include misplacements in PRCL, SRCL and MICH.

 

 



 
 
armsensMAT.png
Figure.2  A sensing matrix of the 40-m DRFPMI while changing the position of ETMX/Y by \sigma = 2 cm.
For convenience,  only REFL11, AS55, POP11 and POP55 are shown. They are the designed signal ports that
mentioned in the aLIGO LSC document (T1000298). In all the histograms, x-axis represents the optical gain in log scale in units of [W/m].
The y-axis is the number of events. The diagonal ports are surrounded by red rectangular window.
 
 
 
 
(Results2 : demodulation phase of MICH and SRCL on POP55)
Now a special attention should be payed on the MICH and SRCL signals on POP55.
Since MICH and SRCL are designed to be taken from POP55, they must be nicely separated in their demodulation phases.
Therefore the demodulation phase of MICH and SRCL has to be carefully examined.
The plot in Figure.3 is the resultant phase difference between MICH and SRCL on POP55 when \sigma_x = \sigma_y = 2 cm.
As shown in the plot the phase are always within a range of 60 - 120 deg, which satisfies my requirement (2) mentioned in the last section.
 
 
 
 
POP55phase2cm.png
 Figure.3 Difference in the demodulation phase of MICH and SRCL on POP55.
x-axis is the difference in the demodulation phase of MICH and SRCL, and y-axis the number of events.
 
 

 

 

  5272   Fri Aug 19 23:41:20 2011 JennyUpdatePSLRelocking NPRO to reference cavity.

Quote:

I am trying again to measure a temperature step response on the reference cavity on the PSL table.

I have been working to relock the NPRO to the cavity. I unblocked the laser beam, reassembled the setup described in my previous elog entry: 5202. I then did the following:

1) Monitored error signal (from LB1005 PDH servo), transmitted signal, and control signal sent to drive PZT on oscilloscope.

2) With loop open, swept through 0,0-mode resonance, saw a peak in the transmission, saw an accompanying error signal similar to the signal shown in 5202.

3) Tried to lock. Swept the gain on the LB1005 and could not find a gain that would make it lock. Tried changing the PI-corner freq. from 10 kHz to 30 kHz and back and still could not lock.

4) Noticed that the open loop error signal displayed on the scope was DC-offset from zero. Changed the offset to zero the error signal.

5) Tried to lock again and succeeded.

6) Noticed that upon closing the loop, the error signal became offset from zero again. Turning on the integrator on the LB1005 increased DC-offset.

7) Reduced the gain on the SR560 being used as a low pass filter from 5 to 1. Readjusted the open loop error signal offset on the LB1005.

8) Closed the loop and locked. Closing the loop then caused a much smaller DC change in the signal than I had seen with the larger gain (now around 3mV). RMS fluctuations in error signal are now 1 mV (well within the linear region of the error signal).

9) Noticed transmission has a strange distorted harmonic oscillation in it a 1MHz. (Modulation freq is 230kHz, so it's not that). Checked reflected signal and also saw a strange oscillation--in a sawtooth-like pattern.

 

I intend to

1) Post oscilloscope traces here showing transmitted and reflected signal when locked.

2) Look upstream to see if the sawtooth-like oscillation is in the laser beam before it enters the cavity:

  • Sweep the temperature of the laser so that the beam is no longer resonating in the cavity.
  • Compare the reflected signal off the cavity to the signal detected before being directed into the cavity (using the PDA255 that I used for measuring the AM response of the PZT) both with and and without the frequency modulation.

3) At some point, try to close the slow digital loop, perhaps readjusting the gain.

4) Try to measure a temperature step response.

I decided to go forward and try to close the digital loop in spite of the unexplained oscillations in the transmission.

1) Plugged the 20dB attenuator into the slow port on the laser drive. This pushed the laser out of lock and, for some reason, made the laser temperature stop responding to sweeping the set point manually with the knob.

2) Plugged the output from the digital system into the slow port (with the attenuator still in place).

3) Displayed the beam seen by the camera on a monitor in the control room

4) Stepped the laser temperature using MEDM until finding the 0,1 mode. Locked to that mode.

5) Closed the digital loop (input to slow laser drive attenuated 20dB attenuator). Gain 0.010

6) Loop appeared stable for 30 minutes, then temperature began shooting off. I opened the loop, cleared history, reduced the gain to 0.008, and started it again. Loop appears stable after 15 minutes of watching. I'm going to leave it for a few hours, then come back to check on it and, if it's stable, step the can temperature.

  5271   Fri Aug 19 19:08:40 2011 JennyUpdatePSLRelocking NPRO to reference cavity.

I am trying again to measure a temperature step response on the reference cavity on the PSL table.

I have been working to relock the NPRO to the cavity. I unblocked the laser beam, reassembled the setup described in my previous elog entry: 5202. I then did the following:

1) Monitored error signal (from LB1005 PDH servo), transmitted signal, and control signal sent to drive PZT on oscilloscope.

2) With loop open, swept through 0,0-mode resonance, saw a peak in the transmission, saw an accompanying error signal similar to the signal shown in 5202.

3) Tried to lock. Swept the gain on the LB1005 and could not find a gain that would make it lock. Tried changing the PI-corner freq. from 10 kHz to 30 kHz and back and still could not lock.

4) Noticed that the open loop error signal displayed on the scope was DC-offset from zero. Changed the offset to zero the error signal.

5) Tried to lock again and succeeded.

6) Noticed that upon closing the loop, the error signal became offset from zero again. Turning on the integrator on the LB1005 increased DC-offset.

7) Reduced the gain on the SR560 being used as a low pass filter from 5 to 1. Readjusted the open loop error signal offset on the LB1005.

8) Closed the loop and locked. Closing the loop then caused a much smaller DC change in the signal than I had seen with the larger gain (now around 3mV). RMS fluctuations in error signal are now 1 mV (well within the linear region of the error signal).

9) Noticed transmission has a strange distorted harmonic oscillation in it a 1MHz. (Modulation freq is 230kHz, so it's not that). Checked reflected signal and also saw a strange oscillation--in a sawtooth-like pattern.

 

I intend to

1) Post oscilloscope traces here showing transmitted and reflected signal when locked.

2) Look upstream to see if the sawtooth-like oscillation is in the laser beam before it enters the cavity:

  • Sweep the temperature of the laser so that the beam is no longer resonating in the cavity.
  • Compare the reflected signal off the cavity to the signal detected before being directed into the cavity (using the PDA255 that I used for measuring the AM response of the PZT) both with and and without the frequency modulation.

3) At some point, try to close the slow digital loop, perhaps readjusting the gain.

4) Try to measure a temperature step response.

  5270   Fri Aug 19 15:31:53 2011 steveUpdateGeneralpower interruption rescheduled to 10-1-2011

                UTILITY & SERVICE INTERRUPTION

**PLEASE POST**

 

Building:               Central Engineering Services (C.E.S.)

          LIGO Gravitational Physics building adjacent to C.E.S. 40M- Lab

          Safety Storage adjacent to CES

          Steele House 

          Keck Lab

 

Date:                   Saturday, October 1, 2011

Time:                   8:00 a.m. To 9:00 a.m.            

Interruption:   Electricity

Contact:                Mike Anchondo ext. 4999  Tom Brennan 4984

*This interruption is required for maintenance of high voltage switchgear in Campus Sub Station.

(If there is a problem with this Interruption, please notify

 the Service Center X-4717 or the above Contact as soon as possible.

 If no response is received we will proceed with the interruption.)

         

                                Reza Ohadi,

                                Director, Campus Operations & Maintenance

  5269   Fri Aug 19 10:26:53 2011 steveUpdateSUSOSEM sensor spectra

Free swingging OSEM sensors LL at atm

Attachment 1: 8freeSUSosemSensors.jpg
8freeSUSosemSensors.jpg
  5268   Fri Aug 19 09:14:56 2011 steveUpdateIOOdegrading laser power at atm

Light into the MC is 20 mW at atm, MC_Transmitted ~10 MW = 400 count

The PMC_T is OK but something else is drifting.

Attachment 1: power@atm.jpg
power@atm.jpg
  5267   Fri Aug 19 01:46:06 2011 SureshUpdateGeneralIFO alignment

[Keiko, Jamie, Kiwamu, Anamaria,

We followed the procedure that we laid-out in our elog of yesterday.  We completed the first six steps and we now have the y-arm well aligned to the green beam which passes through the center of of both ETMY and ITMY. 

The IR beam was steered with the PZTs to coincide with the green beam.  The BS was adjusted to see IR beam scatter on a target placed near the center of the ETMX.  And then the AS IR beam was steered to the AS camera by adjusting several components along OM path ( we touched OM1, OM2, OM3, OM4, OM5, OMPO and OM6).  We then looked for IR fringes in the AS port from the Y-arm. But no luck there.  We need to realign the IR beam into the Y-arm cavity axis using the pzts.

We aligned ITMX and PRM to get power recycled Michelson fringes at the AS.

 

  5266   Fri Aug 19 01:15:22 2011 SureshUpdateSUSFreeSwing all optics

I ran "freeswing all" at Fri Aug 19 01:09:28 PDT 2011 (997776583)  and "opticshutdown"  as well.

 

  5265   Thu Aug 18 22:24:08 2011 jamieOmnistructureVIDEOUpdated 'videoswitch' script

I have updated the 'videoswitch' program that controls the video MUX.  It now includes the ability to query the video mux for the channel mapping:

controls@pianosa:~ 0$ /opt/rtcds/caltech/c1/scripts/general/videoswitch -h
Usage:
videoswitch [options] [OUT]      List current output/input mapping [for OUT]
videoswitch [options] OUT IN     Set output OUT to be input IN

Options:
  -h, --help            show this help message and exit
  -i, --inputs          List input channels and exit
  -o, --outputs         List output channels and exit
  -l, --list            List all input and output channels and exit
  -H HOST, --host=HOST  IP address/Host name
controls@pianosa:~ 0$

  5264   Thu Aug 18 15:54:35 2011 steveUpdateSUSdamped and undamped OSEMs

damped sus at atm1 and freeswingging sus at atm2

 

Attachment 1: 5susLL.jpg
5susLL.jpg
Attachment 2: 8freeSUSLL.jpg
8freeSUSLL.jpg
  5263   Thu Aug 18 12:22:37 2011 jamieUpdateSUSsuspension update

Most of the suspension look ok, with "badness" levels between 4 and 5.  I'm just posting the ones that look slightly less ideal below.

  • PRM, SRM, and BS in particular show a lot of little peaks that look like some sort of intermodulations.
  • ITMY has a lot of elements with imaginary components
  • The ETMY POS and SIDE modes are *very* close together, which is severely adversely affecting the diagonalization
 PRM  PRM.png        pit     yaw     pos     side    butt
UL    0.466   1.420   1.795  -0.322   0.866  
UR    1.383  -0.580   0.516  -0.046  -0.861  
LR   -0.617  -0.978   0.205   0.011   0.867  
LL   -1.534   1.022   1.484  -0.265  -1.407  
SD    0.846  -0.632  -0.651   1.000   0.555

5.62863

SRM SRM.png       pit     yaw     pos     side    butt
UL    0.783   1.046   1.115  -0.149   1.029 
UR    1.042  -0.954   1.109  -0.060  -1.051 
LR   -0.958  -0.926   0.885  -0.035   0.856 
LL   -1.217   1.074   0.891  -0.125  -1.063 
SD    0.242   0.052   1.544   1.000   0.029 
4.0198
 BS BS.png        pit     yaw     pos     side    butt
UL    1.536   0.714   0.371   0.283   1.042  
UR    0.225  -1.286   1.715  -0.084  -0.927  
LR   -1.775  -0.286   1.629  -0.117   0.960  
LL   -0.464   1.714   0.285   0.250  -1.070  
SD    0.705   0.299  -3.239   1.000   0.023 
 5.5501
 ITMY  ITMY.png        pit     yaw     pos     side    butt
UL    1.335   0.209   1.232  -0.071   0.976  
UR   -0.537   1.732   0.940  -0.025  -1.068  
LR   -2.000  -0.268   0.768   0.004   1.046  
LL   -0.129  -1.791   1.060  -0.043  -0.911  
SD   -0.069  -0.885   1.196   1.000   0.239 
 4.28384
ETMY ETMY.png       pit     yaw     pos     side    butt
UL    1.103   0.286   1.194  -0.039   0.994 
UR   -0.196  -1.643  -0.806  -0.466  -1.113 
LR   -2.000   0.071  -0.373  -0.209   0.744 
LL   -0.701   2.000   1.627   0.217  -1.149 
SD    0.105  -1.007   3.893   1.000   0.290 
9.25346

 

Attachment 2: SRM.png
SRM.png
  5262   Thu Aug 18 10:59:04 2011 steveUpdateTreasurePrepared for drag wiping

Quote:

While waiting for the IFO team to align things (there were already ~5 people working on a ~1 person job...), I got all of our supplies prepped for drag wiping in the morning. 

The syringes are still on the flow bench down the Xarm.  I put fresh alcohol from unopened spectrometer-grade bottles into our alcohol drag wiping bottles.

The ITMs already had rails for marking their position in place from the last time we drag wiped.  I placed marker-rails for both ETMs.

 We should use the deionizer before drag wiping with isopropanol.

  5261   Thu Aug 18 10:17:04 2011 kiwamu, steveUpdateSUSoplevs reestablished at Vertex

Kiwamu and Steve, from yesterday

PRM and BS oplev paths were relaid after setting 1/2 OSEM voltages. The incident beam on suspended optics are centered  within  ~ +- 2 mm

I noticed many unvected ss screws are used on the big Al table tops. The SS 1/4-20 screws

 used on the optical tables in vacuum MUST be VENTED!

Also, please use  SS clamps. Replace aluminum ones when you can. We have plenty baked ones.

 

 

  5260   Thu Aug 18 00:58:40 2011 jamieUpdateSUSoptics kicked and left free swinging

ALL optics (including MC) were kicked and left free swinging at:

997689421

The "opticshutdown" script was also run, which should turn the watchdogs back on in 5 hours (at 6am).

 

  5259   Thu Aug 18 00:53:48 2011 jamie, kiwamu, suresh, jenneUpdateGeneralPUMP PLAN ABORTED; need to work more on IFO alignment

We have decided that the IFO alignment is bad enough that we're not ready to pump down.  PUMP ABORTED.

The IFO alignment is somewhat OK, in that the green and IR beams are flashing in the arms, and the return beams are overlapping at the BS.  However the beams appear to be not centered on any of the optics at the moment.  They are all displaced in yaw by ~0.5 to 1 cm or so in various directions.

From this we have decided that we need to step back and reattack the IFO alignment from square one.  Here is our current suggested procedure:

  1. check ETM positions relative to what we think they should be on the drawings.  This is to verify that the ETMs were not placed in the wrong places laterally.
  2. translate Y green axis north, centering green on ETMY and ITMY (by looking at cards).  North is the opposite direction from how the beams are displaced from the TM centers.
  3. steer input pointing to overlap IR on green beam at BS, ITMY, and ETMY.  IR should visibly overlap green at both BS and ITMY, and we should be able to see IR on target in front of ETMY with ETMY face camera, and in ETMY trans camera.
  4. center IR on ETMX by steering BS with DC bias.
  5. align Y arm cavity for green resonance by adjusting ITMY.
  6. adjust ITMX to achieve michelson fringes at AS
  7. adjust PRM lateral translation to center beam on PRM, if needed
  8. adjust SRM lateral translation to center beam on SRM, if needed
  9. align PRC to see fringes
  10. align SRC to see fringes
  11. extract AS (no clipping)

 Once this is done, we will need to check the following:

  • IPANG/IPPOS extraction
  • pick-off extraction
  • OPLEVs
  • OSEMs
  • green periscopes and green beam extraction at PSL

We've decided to stop for the night, get a good nights rest, and attack all of this tomorrow morning.

Beam_spot_shifts.png

  5258   Wed Aug 17 20:14:49 2011 jamie, kiwamu, suresh, jenne, keiko, anamariaUpdateGeneralin-vacuum work status, prep for pump

This afternoon's work:

  • OSEMs were adjusted on all suspended optics.
  • X and Y-arms were aligned to green.
  • Once that was done, the input pointing was adjusted with the PZTs to get the IR beam centered on ITMY and ETMY.
  • Once the input pointing was ok we extracted IPANG and IPPOS.
  • BS, ITMY, PRM, SRM optical levers (oplevs) were extracted.
  • Prepared rails and stops for TMs for morning drag wiping.

TODO before drag wiping:

  • Check full IFO alignment.
  • Readjust OSEMs if needed.
  • Extract ITMX oplev.
  5257   Wed Aug 17 17:51:54 2011 JenneUpdateTreasurePrepared for drag wiping

While waiting for the IFO team to align things (there were already ~5 people working on a ~1 person job...), I got all of our supplies prepped for drag wiping in the morning. 

The syringes are still on the flow bench down the Xarm.  I put fresh alcohol from unopened spectrometer-grade bottles into our alcohol drag wiping bottles.

The ITMs already had rails for marking their position in place from the last time we drag wiped.  I placed marker-rails for both ETMs.

  5256   Wed Aug 17 15:55:01 2011 JenneUpdateGeneralin-vac work : the end is near

Quote:

We will pump down the chambers on Thursday Friday morning.

 All hands on deck at 9am Thursday for drag wiping and doors.  We'll do the 5 doors first (including drag wiping), then put on the access connector last.  Steve will then begin pumping early Friday morning.

ELOG V3.1.3-